5) If Y-A × N × (75 + K/N), where K = 1000, N = 20, and A-10, what happens if K doubles and N doubles? A) Y is unchanged. B) Y increases by 50%. C) Y doubles. D) Y quadruples.show work
Answer #1

Y = AN(75 + KN)

Y = 10 x 20 (75 + 1000 / 20)

Y = 200 (75 + 50)

Y = 15000 + 10,000

Y = 25,000.

If K and N doubles then Y = 10 x 40 ( 75 + 2000 / 40).

400 (75 + 50)

= 30,000 + 20,000

= 50,000.

the answer is "C" Y doubles.